During 2017, Columbia Inc. had beginning accounts receivable of $680,000 and ending accounts receivable of $760,000. Its net sales of $4,500,000 are composed of 20% cash sales and 80% credit sales. Based on this information, what is Columbia's average collection period?

Group of answer choices

58.4 days

292.0 days

73.0 days

5.0 days

None of the answer choices is correct.
Answer #1

Accounts Receivable Turnover = Credit sales/Average accounts receivables

= (4,500,000*80%) / [(680,000+760,000)/2]

= 3,600,000/720,000

= 5 times

Average collection period = 365/Accounts Receivable Turnover l

= 365/5

= 73 days